# Persuasion in Networks: Can the Sender Do Better than Using Public Signals?

### v2

@@ Abstract
Political and advertising campaigns increasingly exploit social networks to spread information and persuade people. This paper studies a persuasion model to examine whether such a strategy is better than simply sending public signals. Receivers in the model have heterogeneous priors and will pass on a signal if they are persuaded by it to take sender’s preferred action. I show that a risk neutral or risk loving sender prefers to use public signals, unless more skeptical receivers are sufficiently more connected in the social network. A risk averse sender may prefer to exploit the network. These results still hold when the network exhibits homophily.
@@
To see the trade-off between using mass media and social networks, consider an example where a pharmaceutical firm tries to advertise its new drug. Some customers are skeptical about the drug while others have prior belief that the drug is safe and effective. The firm would like to send the result of a clinical trial to the skeptics, but keep such information away from the believers, since negative results from the trials could dissuade the believers. If mass media is used to disseminate information, everyone must receive the same information. With information disseminated on social networks, the firm could try to target information towards the sceptics. However, people will also be sharing information with each other, so it is unclear whether and when the firm can succeed in sending information more often to sceptics than believers, and thus unclear whether the firm prefers public dissemination or dissemination using social network. In addition, the risk properties of the two options are different. Everyone with the same prior must choose the same action with public dissemination, whereas people can have different observations and thus use different actions when information is spread on a network.
To address these questions, I propose the following persuasion model. There are two states and receivers have two actions $\{0,1\}$. There are infinitely many receivers, who want to match their action with the state. The sender wants receivers to take action 1 regardless of the state. Receivers have heterogeneous priors, some of them (believers) would choose action 1 while others (sceptics) would choose action 0 if no further information is provided. Receivers are located on a network captured by a random network model. Receivers know their own degree.
@@
I allow the sender to send no signal. This plays no role with public signals. If sender uses network signals, receivers may receive no information due to the network structure.When receivers observe no signal, they cannot distinguish whether it is due to network structure or sender choosing to send no signal. The sender could potentially exploit this in her benefit.
The main question in this paper is whether and when network signals can outperform public signals for the sender. The potential cost of network signals is that some sceptics may not receive information due to structure of the network and/or how signals are shared. The potential benefit of network signal is that the sender may be able to send information more often to the sceptics than to believers. Also, as noted above, network and public signals have different risks. Conditions under which the potential benefit can and cannot be achieved are identified. This in turn allows us to compare public and network signals, taking into account sender’s risk preferences.
@@
The first main result states that in the baseline network (i.e. one without homophily or correlations between degree and priors), if the sender is risk loving or risk neutral, network signals cannot outperform public signals. It is possible for network signals to perform strictly better if the sender is risk averse. Believers and skeptics will be ex ante symmetric in terms of their positions in such networks. Therefore, there are no differences between how skeptics and believers are exposed to signals, and the sender cannot send information more often to sceptics than believers. The sender can only prefer network signals due to its risk properties.
Network signals in general result in receivers sometimes not observing any signal. With public signals, if receivers do not receive good news, they know with certainty that good news is not realised. With network signals, the event of not observing good news is less informative since it may be due to the network structure. This allows the sender to increase the probability of sending good news, without dissuading believers when good news is not realised. This comes at the cost that when good news does realise, less receivers observe it. Therefore, network signals allow the sender to reduce risk, so could be beneficial if she is risk averse.
@@
Introducing homophily or making believers more connected than sceptics does not change these results. Networks exhibiting these features do introduce asymmetry between believers and sceptics. However, the asymmetry will actually make believers more exposed to signals than sceptics, which is the opposite of what the sender desires.
@@
If skeptics are sufficiently more connected in the network than believers, the sender prefers network signals regardless of her risk preferences. Skeptics being more connected allows the sender to send information to them more often than to the believers.
@@
In summary, this paper compares performance of public and network signals. In particular, it shows that under many plausible conditions, public signals perform better. Network signals can be preferred by the sender only when the sender is risk averse and/or sceptics are more connected. In practice, sender in many situations have access to tools for public dissemination of information. The results imply that, in many situations, resources may be more efficiently spent on sending public signals, rather than on trying to understand network structures and picking the optimal seeds. In addition, persuasion on large networks is complicated while public signals have been well studied. From these comparisons, we can use sender’s value from public signals to obtain what sender could get at most or at least using network signals, whose value for the sender can be very difficult to compute.
@@ 1 Literature Review
This paper is most related to the literature on persuasion with information spillovers. Galperti and Perego (2023) studies an environment where a sender can directly send signals to some receivers (seeds), and receivers observe the same information as those connected to them (by a directed path). They show that the problem is equivalent to allowing the sender to directly send information to all receivers, by transforming the network appropriately. They provide a characterization of feasible outcomes and show how it changes with the network and seeds. Babichenko et al. (2021) studies a similar problem, and also explores the algorithmic approach. Kerman and Tenev (2021) studies a voting game in which receivers observe their neighbours’ signals. They characterise sender’s optimal signals in some stylised networks, and compare sender’s payoff to the case when receivers do not communicate. Liporace (2021) assumes that receivers observe signals of their neighbours, and receivers have heterogeneous priors. She characerises sender’s optimal strategy, restricting attention to strategies that persuade everyone in the good state. In Egorov and Sonin (2020), information is shared between receivers with some probability, and receivers choose whether to subscribe to the sender which is costly. They characterise sender’s optimal strategy in some stylised networks.
@@
This strand of literature usually assumes that information will be transmitted mechanically to one’s neighbours, while this paper studies more endogenous spread of information. Another key difference is that, the sender in these papers can always make sure that all receivers observe information from the sender, so they do not consider the possibility that sender’s information may not reach some receivers, which is very likely when considering large social networks. Also, in this paper, sender and receivers do not have full information of the network, whereas the literature studies fixed and known networks. Uncertainties over network is a plausible assumption in large networks.
@@
Also related is the literature on strategic communication, which studies how players communicate when their payoffs depend on each other’s actions. Most papers in this literature (e.g. Galeotti et al. (2013) and Hagenbach and Koessler (2010)) focus on conditions under which players truthfully communicate with each other, and do not study persuasion problems. The exception is Squintani (2024). Squintani (2024) studies an environment where there is one sender and one receiver (randomly selected from all players). He characterises conditions for truthful communication and studies the optimal network and network formation.
@@
Some papers have looked at the comparison between public signals and private signals, but do not allow communication between receivers. Wang (2013) compares public signals and i.i.d. signals in a voting game, and finds that public signal is always better. Inostroza and Pavan (2018) also compares public and private signals, imposing certain restrictions on private signals. Candogan (2020) studies a problem of a social media platform to signal the quality of content, and finds that public signals are sufficient if the goal is to minimize misinformation. Arieli and Babichenko (2019) relates optimality of public signals to the core of a cooperative game. Communication between receivers puts endogenous restrictions on joint distributions over what receivers observe when sender sends personalised signals. Also, receivers have heterogeneous priors in this paper. This potentially makes personalised signals more attractive. Due to these and other more detailed differnces in setup, the trade-offs and insights in this paper are different from existing papers in this literature.
@@
Finally, this paper uses random networks to model the social network. Random networks have been used in Akbarpour et al. (2020), Sadler (2020), Campbell (2013) and Golub and Jackson (2012) among others, to study various problems.
@@ 2.1 Agents and Network
I consider an environment with one sender, and an infinite and countable set of receivers. There are two possible states of the world, $\omega\in\{0,1\}$, and each receiver has two possible actions $\{0,1\}$. The receivers want to match their action with the state. The receiver gets a payoff of 1 if their action matches the state, and gets 0 otherwise. The sender wants receivers to choose action 1 regardless of the state. If a proportion $x$ of the receivers choose action 1, then the sender gets payoff $v(x)$, where $v(.)$ is continuous, weakly increasing and bounded. I normalise $v(0)=0$.
@@
A proportion $\gamma_{h}$ of the receivers are of type $h$ and have prior belief that state $\omega$ realises with probability $\mu_{h}(\omega)$. The rest of the receivers are of type $l$, and have a prior $\mu_{l}(\omega)$ for state $\omega$. I assume that $\mu_{h}(\omega=1)>0.5>\mu_{l}(\omega=1)$. I denote the proportion of type $l$ receivers by $\gamma_{l}$. Type $h$ receivers are often referred to as believers, and type $l$ receivers as sceptics.
@@
The sender has a prior belief that state $\omega$ realises with probability $\mu_{s}(\omega)$. The sender can send signals from the signal space $S\cup\{\emptyset\}$, where $S$ is finite and contains at least two elements [^5] All results below easily extend to cases where $S$ has only one element or $S$ is empty. If $S$ is empty, the result is trivial since the sender essentially cannot do anything. If $S$ contains one element, payoff from network signals weakly decreases, while it can be easily shown that payoff using public signals does not change. For results where network signals perform better than public signals, in their proofs, sender only uses 1 signal from $S$ with positive probability.. The element $\emptyset$ captures the sender’s ability to send no signal. The sender designs a signal structure $\pi(.)$. The probability of signal $s\in S\cup\{\emptyset\}$ being realised in state $\omega$ is $\pi(s|\omega)$, and $\pi(.)$ must satisfy $\sum_{s\in S\cup\{\emptyset\}}\pi(s|\omega)=1$ for $\omega=0,1$. If the sender uses network signals, the sender can also choose where to seed the signals. Precisely how seeding works and timing of the game are described in the next two subsections.
@@
Now we describe the random network model, which captures how receivers are connected to each other. Each receiver $i$ has connectedness $\lambda_{i}$. With $n$ receivers, if $i$ and $j$ are have the same prior, the probability that they are connected is $min\{\frac{\lambda_{i}\lambda_{j}}{n},1\}$. If they have different priors, the probability that they are connected is $min\{\frac{q\lambda_{i}\lambda_{j}}{n},1\}$ where $q\in(0,1]$. Intuitively, $q$ controls homophily. Whether $i$ and $j$ are connected is independent of connections between all other pairs (conditional on their connectedness and types). Conditional on having type $t$, a receiver has connectedness $\lambda$ with probability $f_{t}(\lambda)$, where $f_{t}(\cdot)$ is a probability mass function with finite support [^6] The analysis can be extended to the case where distribution over $\lambda$ is continuous and unbounded. This will make restrictions on these distributions and the proofs more tedious, without adding much insight. Bollobás et al. (2007) states the precise restrictions on these distributions when they have continuous and unbounded support..
@@
Note that the probability that two nodes are connected depends on the number of receivers $n$. There are infinitely many receivers in the model, so we look at the limiting network when $n$ goes to infinity, as in many other papers using random network models. The network formation process is common knowledge. Each receiver $i$ in addition knows her own degree, denoted by $d_{i}$ (degree is the number of neighbours a receiver has). When studying network signals, I use receivers and nodes exchangeably.
@@
Two important network statistics are degree distribution and the probability that a node is connected to a node of the same type. Theorem 3.13 in Bollobás et al. (2007) shows that
@@
As $n\rightarrow\infty$, degree distribution of type $t$ receivers converges to $\{p_{d}^{t}\}_{d}$. Also, conditional on degree $d$ and type $t$ of a node, each of its neighbour has an independent probability $q_{t}$ of having the same type. The degree distribution $\{p_{d}^{t}\}_{d}$ and probability $q_{t}$ are
@@ 2.2 Network signals
As described above, social network with infinitely many receivers is modelled as the limit of a sequence of finite random networks. Therefore, I will first describe the game with $n$ receivers. Sender’s optimal payoff facing infinitely many receivers will be captured as the limit of her optimal payoffs in these finite games.
With $n$ receivers, the timing goes as follows. The sender moves first and chooses a signal structure $\pi(.)$, and a seeding strategy $Z(.)$. Given a network with $n$ receivers $g^{n}$ and a signal $s$, $Z(g^{n},s)$ returns a subset of receivers. The subset selected by $Z(.)$ must have a smaller size than $n^{\alpha}$, where $\alpha<1$. Intuitively, $Z(.)$ selects some seeds that will receive signals first, and seed selection can be conditioned on the network and the signal realised.
After seeing the sender’s strategy, the receivers choose their actions for each degree $d$ and observation in $S\cup\{\emptyset\}$, $a_{i}:(S\cup\emptyset)\times\mathbb{N}\rightarrow\{0,1\}$, where $a_{i}(s,d)$ describes what $i$ would do when she has degree $d$ and observes $s\in S\cup\{\emptyset\}$.
Then nature draws a network $g^{n}$ according to the network formation process described above. Nature draws a state, a signal $s$ according to $\pi(.)$. If the signal is $\emptyset$, all receivers observe $\emptyset$. Intuitively, if the sender sends out no signal, then all receivers observe no signal.
If $s\neq\emptyset$, then seeds are selected according to $Z(.)$. Given network $g^{n}$ and the seeds, a receiver $i$ observes $s$ if and only if at least one of the following conditions hold:
@@
All the other receivers observe $\emptyset$. In words, the condition here says that a receiver will pass on a signal to her neighbours, if and only if she is persuaded by the signal to take sender’s preferred action. Intuitively, people will pass on the advertisement of a product or a politician if they are convinced to buy the product or vote for the politician.
The solution concept is subgame perfect equilibrium. Note that, although we assume that the receivers move before realisation of their degree and signals, we are allowing their actions to be conditioned upon these information. Therefore, it is essentially the same as assuming that they move after observing their degrees and the signal. This is standard in Bayesian Persuasion (Kamenica and Gentzkow, 2011) and when studying Bayesian games more generally.
For each $n$, we can find the sender’s maximum (or supremum) equilibrium payoff. Denote this payoff by $V^{*}_{net}(n)$. Sender’s payoff facing infinitely many receivers is captured by $lim_{n\rightarrow\infty}V^{*}_{net}(n)$. However, this limit may not exist for some parameters [^7] It can be shown that this only happens for a non-generic set of parameters, but the proof is very long and does not bring additional insight, so I will not discuss it for the sake of space.. Therefore, I compare $limsup_{n\rightarrow\infty}V^{*}_{net}(n)$ and $liminf_{n\rightarrow\infty}V^{*}_{net}(n)$ with the sender’s payoff from public signals.

### v3

@@ Abstract
Political and advertising campaigns increasingly exploit social networks to spread information and persuade people. This paper studies a persuasion model to examine whether such a strategy is better than simply sending public signals. Receivers in the model have heterogeneous priors and will pass on a signal if they are persuaded by it to take sender’s preferred action. I show that a risk-neutral or risk-loving sender prefers to use public signals, unless the more sceptical receivers are sufficiently more connected in the network. This result still holds when the network exhibits homophily. When the sender is risk averse or has a threshold function as payoff function, I characterise conditions under which the sender prefers to exploit the network. An important factor identified by these conditions is whether the less sceptical receivers are sufficiently connected.
@@
To see the trade-off between using mass media and social networks, consider an example in which a firm tries to advertise its product. Some consumers are sceptical about the quality of the product, while others have prior belief that it has high quality. In order to persuade the sceptics to buy, the firm can conduct quality tests or get pre-release reviews, and send sceptics this information. Ideally, the firm would like to send this information to the sceptics but keep them away from the believers, since negative results from the tests and negative reviews could dissuade the believers. If the information is broadcast publicly, everyone must receive the same information. If the firm spreads information on social networks, some people may not receive information due to structure of the network and the way people share information, and the firm may be able to exploit this. In particular, the firm could try to target information towards the sceptics. However, communication over the network endogenously imposes restrictions on correlations between what believers and sceptics observe, so it is unclear whether and when the firm can succeed in sending information more often to sceptics than to believers.
In this setup, does the firm want to publicly broadcast information or spread information on the network? What are the networks in which public broadcasts can and cannot be outperformed? How do the firm’s preferences influence the comparison?
To address these questions, I propose the following persuasion model. There are two states and receivers have two actions $\{0,1\}$. There are infinitely many receivers, who want to match their action with the state. The sender wants receivers to take action 1 regardless of the state. Receivers have heterogeneous priors. If receivers make decisions based on priors, some of them (believers) would choose action 1 while others (sceptics) would choose action 0. Receivers are located on a network captured by a random network model. Receivers know their own degree.
@@
I allow the sender to send no signal. This plays no role with public signals. If the sender uses network signals, the receivers may not receive any signals due to the network structure. When receivers observe no signal, they cannot distinguish whether it is due to network structure or sender choosing to send no signal. The sender could potentially exploit this in her benefit.
The main research question is whether and when network signals can outperform public signals for the sender. The answer to this question in turn depends on whether the sender can use network signals and seeding to send information more often to sceptics than to believers. Also, some sceptics may not receive any signals if the sender uses network signals. This may be beneficial or costly to the sender, depending on the sender’s preference.
@@
The first result states that in the baseline network (i.e. one without homophily or correlations between degree and priors), if the sender is risk loving or risk neutral, network signals cannot outperform public signals. The main intuition underlying this result is that the sender needs signals to get viral in order to reach a large population, but it is difficult to control who are reached by a viral signal using seeding. Also, believers and sceptics will be ex ante symmetric in terms of their positions in such baseline networks. Therefore, there are no differences between how sceptics and believers are exposed to signals, and the sender cannot send information more often to sceptics than to believers.
@@
One might suspect that homophily allows the sender to exploit seeding and provide information more often to sceptics than to believers. However, introducing homophily and/or making believers more connected than sceptics does not change the above result. Networks exhibiting these features do introduce asymmetry between believers and sceptics. However, the asymmetry makes believers more exposed to signals than sceptics, which is the opposite of what the sender desires.
@@
If sceptics are sufficiently more connected in the network than believers, the sender prefers network signals. Sceptics being more connected allows the sender to send information to them more often than to the believers.
Next, I study the case where the sender is risk averse. When believers are sufficiently connected in the network, the sender can always persuade some sceptics using network signals, but it is in general impossible to persuade all sceptics at the same time. The reverse is true for public signals. Therefore, when believers are sufficiently connected, a sufficiently risk-averse sender prefers network signals.
Finally, I study a voting game, in which sender’s payoff function is a threshold function. Like in the case of a risk-averse sender, it is found that whether believers are sufficiently connected is an important factor influencing whether the sender prefers network signals.
In summary, this paper characterises conditions on the network structure and the sender’s preference under which network signals can and cannot outperform public signals. It shows that in many cases, public signals perform better even if receivers have heterogeneous priors. The conditions also illustrate that the sender’s preference will influence which aspect of the network matters for the comparison between public signals and network signals.
@@
In practice, the sender in many situations has access to tools for publicly broadcasting information. The results imply that, in many cases, resources may be more efficiently spent on sending public signals, rather than on trying to understand network structures and picking the optimal seeds. In addition, persuasion on large networks is complicated while public signals have been well studied. The sender’s value of using network signals can be very difficult to compute. From these comparisons, we can use sender’s value of using public signals to obtain what sender could get at most or at least using network signals.
@@ 1 Literature Review
This paper is most related to the literature on persuasion with information spillovers. Galperti and Perego (2023) studies an environment where a sender can directly send signals to some receivers (seeds), and receivers observe the same information as those connected to them (by a directed path). They show that the problem is equivalent to allowing the sender to directly send information to all receivers, by transforming the network appropriately. They provide a characterization of feasible outcomes and show how it changes with the network and seeds. Babichenko et al. (2021) studies a similar problem, and also explores the algorithmic aspect of the problem. Kerman and Tenev (2021) studies a voting game in which receivers observe their neighbours’ signals. They characterise sender’s optimal signals in some stylised networks, and compare sender’s payoff to the case when receivers do not communicate. Liporace (2021) assumes that receivers observe signals of their neighbours, and receivers have heterogeneous priors. She characterises the sender’s optimal strategy, restricting attention to strategies that persuade everyone in the good state. In Egorov and Sonin (2020), information is shared between receivers with some probability, and receivers choose whether to subscribe to the sender which is costly. They characterise the sender’s optimal strategy in some stylised networks.
@@
This strand of literature usually assumes that information will be transmitted mechanically to one’s neighbours, while this paper studies more endogenous spread of information. Another key difference is that, the sender in most of these papers can always make sure that all receivers observe information from the sender. They do not consider the possibility that sender’s information may not reach some receivers, which is very likely when considering large social networks. Also, in this paper, the sender and receivers do not have full information of the network, whereas the literature studies fixed and known networks. Uncertainties over network is a plausible assumption in large networks.
@@
Also related is the literature on strategic communication, which studies how players communicate by cheap talk when their payoffs depend on each other’s actions. Most papers in this literature (e.g. Galeotti et al. (2013) and Hagenbach and Koessler (2010)) focus on conditions under which players truthfully communicate with each other, and do not study persuasion. The exception is Squintani (2024). Squintani (2024) studies an environment where there is one sender and one receiver (randomly selected from all players). He characterises conditions for truthful communication and studies the optimal network and network formation.
@@
Some papers have looked at the comparison between public signals and private signals, but do not allow communication between receivers. Wang (2013) compares public signals and i.i.d. signals in a voting game, and finds that public signal is always better. Inostroza and Pavan (2018) also compares public and private signals, imposing certain restrictions on private signals. Candogan (2020) studies a problem of a social media platform to signal the quality of content, and finds that public signals are sufficient if the goal is to minimize misinformation. Arieli and Babichenko (2019) relates optimality of public signals to the core of a cooperative game. Communication between receivers puts endogenous restrictions on joint distributions over what receivers observe when sender sends personalised signals. Due to this and other more detailed differences in setup, the trade-offs and insights in this paper are different from existing papers in this literature.
@@
Finally, this paper uses a random network model. Random networks have been used in Akbarpour et al. (2020), Sadler (2020) and Campbell (2013) to model social networks among others. In particular, it is very often used to model diffusion of ideas and information, see e.g. Golub and Jackson (2012), Sadler (2023) and Merlino et al. (2023). The problems studied in these papers are very different from this paper.
@@ 2.1 Agents and Network
I consider an environment with one sender, and an infinite and countable set of receivers. There are two possible states of the world, $\omega\in\{0,1\}$, and each receiver has two possible actions $\{0,1\}$. The receivers want to match their action with the state. The receiver gets a payoff of 1 if their action matches the state, and gets 0 otherwise. The sender wants receivers to choose action 1 regardless of the state. If a proportion $x$ of the receivers choose action 1, then the sender gets payoff $v(x)$, where the function $v$ is continuous, weakly increasing and bounded. I normalise $v(0)=0$.
@@
A proportion $\gamma_{h}$ of receivers are of type $h$ and have prior belief that state $\omega$ realises with probability $\mu_{h}(\omega)$. The rest of the receivers are of type $l$, and have a prior belief $\mu_{l}(\omega)$ for state $\omega$. I assume that $\mu_{h}(\omega=1)>0.5>\mu_{l}(\omega=1)$. I denote the proportion of type $l$ receivers by $\gamma_{l}$. Type $h$ receivers are often referred to as believers, and type $l$ receivers as sceptics.
@@
The sender has a prior belief that state $\omega$ realises with probability $\mu_{s}(\omega)$. The sender can send signals from the signal space $S\cup\{\emptyset\}$, where $S$ is finite and contains at least two elements [^5] All results below easily extend to cases where $S$ has only one element or $S$ is empty. If $S$ is empty, the result is trivial since the sender essentially cannot do anything. If $S$ contains one element, payoff from network signals weakly decreases, while it can be easily shown that payoff using public signals does not change. For results where network signals perform better than public signals, in their proofs, sender only uses 1 signal from $S$ with positive probability.. The element $\emptyset$ captures the sender’s ability to send no signal. The sender designs a signal structure $\pi:\{0,1\}\rightarrow\Delta(S\cup\{\emptyset\})$. The probability of signal $s\in S\cup\{\emptyset\}$ being realised in state $\omega$ is $\pi(s|\omega)$, and $\pi$ must satisfy $\sum_{s\in S\cup\{\emptyset\}}\pi(s|\omega)=1$ for $\omega=0,1$. If the sender uses network signals, the sender can also choose where to seed the signals. Precisely how seeding works and timing of the game are described in the next two subsections.
@@
Now we describe the random network model. Each receiver $i$ has connectedness $\lambda_{i}$. With $n$ receivers, if $i$ and $j$ have the same prior, the probability that they are linked is $min\{\frac{\lambda_{i}\lambda_{j}}{n},1\}$. If they have different priors, the probability that they are linked is $min\{\frac{q\lambda_{i}\lambda_{j}}{n},1\}$ where $q\in(0,1]$. Intuitively, $q$ controls homophily. Whether $i$ and $j$ are linked is independent of links between all other pairs (conditional on connectedness of $i$ and $j$). Conditional on having type $t$, a receiver has connectedness $\lambda$ with probability $f_{t}(\lambda)$, where $f_{t}(\cdot)$ is a probability mass function with finite support [^6] The analysis can be extended to the case where distribution over $\lambda$ is continuous and unbounded. This will make restrictions on these distributions and the proofs more tedious, without adding much insight. Bollobás et al. (2007) states the precise restrictions on these distributions when they have continuous and unbounded support..
@@
Note that the probability that two nodes are linked depends on the number of receivers $n$. There are infinitely many receivers in the model, so we look at the limiting network when $n$ goes to infinity, as in many other papers using random network models. The network formation process is common knowledge. Each receiver $i$ in addition knows her own degree, denoted by $d_{i}$ (degree is the number of neighbours a receiver has). When studying network signals, I use receivers and nodes interchangeably.
The random network model is borrowed from Bollobás et al. (2007). As discussed in the literature review, random network models are widely applied to model large social networks. In particular, they have been used to capture networks over which people communicate information and opinions.
@@
Two important network statistics are degree distribution and the probability that a node’s neighbour has the same type as the node. Theorem 3.13 in Bollobás et al. (2007) shows that
@@
As $n\rightarrow\infty$, degree distribution of type $t$ receivers converges to $\{p_{d}^{t}\}_{d}$. Also, conditional on degree $d$ and type $t$ of a node, each of its neighbours has an independent probability $q_{t}$ of having the same type [^7] More precisely, as $n\rightarrow\infty$, conditional on degree $d$ and type $t$ of a node, the probability that $d_{t}$ of its neighbours are of type $t$ converges to $Binom(d_{t};d,q_{t})$, i.e. type $t$. The degree distribution $\{p_{d}^{t}\}_{d}$ and probability $q_{t}$ are
@@ 2.2 Network signals
As described above, social network with infinitely many receivers is modelled as the limit of a sequence of finite random networks. Therefore, I will first describe the game with $n$ receivers. The sender’s optimal payoff facing infinitely many receivers is captured as the limit of her optimal payoffs in these finite games when $n\rightarrow\infty$.
With $n$ receivers, the timing goes as follows. The sender moves first and chooses a signal structure $\pi$, and a seeding strategy $Z$. Given a network with $n$ receivers $g^{n}$ and a signal $s\in S$, $Z(g^{n},s)$ returns a subset of receivers. The subset selected by $Z$ must have a size smaller than $n^{\alpha}$, where $\alpha<1$. Intuitively, $Z$ selects some seeds that will receive signals first, and seed selection can be conditioned on the network and the signal realised.
After seeing the sender’s strategy, the receivers choose their strategies, $a_{i}:(S\cup\emptyset)\times\mathbb{N}_{0}\rightarrow\{0,1\}$, where $a_{i}(s,d)$ describes what $i$ would do when she has degree $d$ and observes $s\in S\cup\{\emptyset\}$.
Then nature draws a network $g^{n}$ according to the network formation process described above. The nature draws a state, a signal $s$ according to $\pi$. If the signal is $\emptyset$, all receivers observe $\emptyset$. Intuitively, if the sender sends no signal, then all receivers observe no signal.
If $s\neq\emptyset$, then the receivers $Z(g^{n},s)$ are selected as seeds. Given network $g^{n}$ and the seeds, a receiver $i$ observes $s$ if and only if at least one of the following conditions hold:
@@
All the other receivers observe $\emptyset$. In words, the condition here says that a receiver will pass on a signal to her neighbours, if and only if she is persuaded by the signal to take the sender’s preferred action. Intuitively, people will pass on the advertisement of a product or a politician if they are convinced to buy the product or vote for the politician.
As in standard Bayesian persuasion, since receivers move after the sender, their strategies must constitute a (Nash) equilibrium in all subgames after the sender moves. Given receivers’ equilibrium behaviour in the subgames, the sender chooses a strategy to maximise her expected payoff. Equilibrium in subgames and the sender’s problem are described in more detail below.
Let $\Pi\equiv(\pi,Z)$ denote the sender’s network signal strategy and $\mathbf{a}_{-i}$ denote the strategy profile of receivers excluding $i$. Given $\Pi$ and $\mathbf{a}_{-i}$, receiver $i$ with type $t_{i}$ form a posterior upon observing each $s\in S\cup\{\emptyset\}$ and her degree $d\in\mathbb{N}_{0}$. The posterior belief for state 1 is denoted by $\psi_{net}(s,d;\mathbf{a}_{-i},\Pi,t_{i})$. Recall that receivers optimally choose action 1 if and only if they believe state 1 happens with probability higher than 0.5, so equilibrium in the subgame following $\Pi$ is defined as
Definition 1.
In a network signals game, receiver $i$’s strategy $a_{i}$ is a best response to $\Pi$ and $\mathbf{a}_{-i}$ if $a_{i}(s,d)=1\iff\psi_{net}(s,d;\mathbf{a}_{-i},\Pi,t_{i})\geq 0.5$ for all $s\in S\cup\{\emptyset\}$ and $d\in\mathbb{N}_{0}$. In the subgame following $\Pi$, receivers’ strategy profile $\mathbf{a}$ is an equilibrium if $a_{i}$ is best responding to $\Pi$ and $\mathbf{a}_{-i}$ for all $i$.
The next remark shows that, given any $\Pi$, the subgame following it has a unique equilibrium outcome. [^8] This remark relies on breaking indifferences in favour of the sender as assumed above. Again, as stated above, results below still go through if we break indifferences against the sender..
Remark 2.
For any sender’s strategy $\Pi$, all equilibria in the subgame following $\Pi$ induce the same distribution over receiver’s action profile in each state.
The remark implies that the sender’s expected payoff is the same in all equilibria in the subgame following $\Pi$. Let $V_{net}(\Pi,n)$ denote sender’s expected payoff in the network signals game with $n$ receivers, when the sender chooses $\Pi$ and receivers play according to equilibrium in the subgame following $\Pi$. The sender’s problem is to choose $\Pi$ and maximise $V_{net}(\Pi,n)$
For the game with $n$ receivers, we can find the sender’s supremum payoff. Denote this payoff by $V^{*}_{net}(n)$. Sender’s payoff facing infinitely many receivers is captured by $lim_{n\rightarrow\infty}V^{*}_{net}(n)$. However, this limit may not exist for some parameters [^9] It can be shown that this only happens for a non-generic set of parameters, but the proof is very long and does not bring additional insight, so it is relegated to online appendix for the sake of space.. Therefore, I compare $limsup_{n\rightarrow\infty}V^{*}_{net}(n)$ and $liminf_{n\rightarrow\infty}V^{*}_{net}(n)$ with the sender’s payoff from public signals.